A random variable X has the following probability function : x: -2 -1 0 1 2 3 f(x): 0*1 k 0*1 2 2k 0*3 3k (i) Determine the distribution function F(x) of X.

1
Expert's answer
2021-11-15T15:59:43-0500
"0.1+k+0.12+2k+0.3+3k=1"

"6k=0.48"

"k=0.08"

"F_X(x) = \begin{cases}
 0 & x<-2 \\
 0.1 &-2\leq x<-1\\
 0.18 &-1\leq x<0\\
 0.3 &0\leq x<1\\
 0.46 &1\leq x<2\\
 0.7 &2\leq x<3\\
 1 &x\geq 3\\
\end{cases}"